Intial testing of BS04 data.
BS04cfg005 = Final assembly in MOBY for the MOBY264 deployment. After sys response data (fri night) where collected the FO block was remove, carted from tent to Rm 214 for Mark to bolt together optics with the ctrlr+pwr+cooling. On Sunday the housed BS was installed in MOBY, and the FO-block was re-attached. Because of the FO block disconnect Mike decided to increase the config number by one.Problems seen (and solved) and good things:
System responses (with new wave cals)
Track Wavecals (day 4 and 5 with 5 pen lamps - 3rd order polynomial) - new on 11 Dec 2018 - still need to redo the rspsP(2,:) = [-5.477399e-09 2.01442e-05 0.3343458 341.1147] Es - Track 2 wavecals from Page number 5.03 P(6,:) = [-5.712727e-09 2.045533e-05 0.3341945 341.1941] LuTop - Track 6 wavecals from Page number 5.01 P(9,:) = [-5.868676e-09 2.065226e-05 0.3342042 341.1773] LuMid - Track 9 wavecals from Page number 5.02 P(13,:) = [-6.972174e-09 2.203839e-05 0.3340448 341.1936] LuBot - Track 13 wavecals from Page number 4.01
Postcal int time cals
All the raw data files and their images , Table of KEYWORDS variable | |||
Page Number |
Link |
Description |
Date |
1.01 |
Day 01 - 23 May Raw Data , Table |
DAY01 - LED FlashLight scans to verify my camera & shutter control which didn't work well on day00 |
5 Jun 2018 |
1.02 |
Day 02 - 25 May Raw Data , Table |
DAY02 - First set of Lu post-cals via the OL425 lamp -
scan sets were 2xDk,4xLt,2xDk via LuBot,Mid,Top |
5 Jun 2018 |
1.03 |
Day 03 -7 Jun Raw Data , Table |
DAY03 -BS04 Es post-cal using the GS5000 cfg F600(c3) |
7 Jun 2018 |
1.04 |
Day 04 -14 Jun Raw Data , Table |
DAY04 - LuBot wave cal for post-cal |
14 Jun 2018 |
1.05 |
Day 05 -15 Jun Raw Data , Table |
DAY05 -5x wave lamps on the rest of the post-M264 BS04 Lu & Es sensors |
14 Jun 2018 |
1.06 |
Day 06 -21 Jun Raw Data , Table |
DAY06 - post-M264 BS04 LuBot exposure time cal |
14 Jun 2018 |
1.07 |
Day 07 -29 Jun Raw Data , Table |
DAY07 - post-M264 BS4 LuBot via Ekspla |
14 Jun 2018 |
1.08 |
Day 08 -17 Jul Raw Data , Table |
DAY08 - post-M264 BS4 LuBot via Ekspla at 410 nm saturated and unsaturated |
26 Jun 2018 |
Day02 - LED FlashLight scans to verify my camera & shutter control which didn't work well on day00 | |||
2.01 | findtracks | Find the full track definition | 28 Aug 2018 |
2.02 | Checking saddle shape and track movement precal, in-water to postcals | 28 Aug2018 | |
2.03 | lu rsp | Postcal Lu rsp | 28 Aug 2018 |
DAY03 -BS04 Es post-cal using the GS5000 cfg F600(c3) | |||
3.01 | es rsp | Postcal Es rsp | 28 Aug 2018 |
DAY04 - LuBot wave cal for post-cal | |||
4.01 | LuBot Wavecal | LuBot 5 pen lamp wavecal | 28 Aug 2018 |
DAY05 - 5x wave lamps on the rest of the post-M264 BS04 Lu & Es sensors | |||
5.01 | LuTop Wavecal | LuTop 5 pen lamp wavecal | 28 Aug 2018 |
5.02 | LuMid Wavecal | LuMid 5 pen lamp wavecal | 28 Aug 2018 |
5.03 | Es Wavecal | Es 5 pen lamp wavecal | 28 Aug 2018 |
5.04 | trying to calc track wavecals | Can you calc other track wavelengths from 4 track pen lamp wavecal | 28 Aug 2018 |
5.05 | Pen lamp line movement by track | Looking at the pen lamp line movement by track number | 28 Aug 2018 |
DAY06 - post-M264 BS04 LuBot exposure time cal | |||
6.01 | int time cals | LuBot int time cals (Brite, Mid, dim and dimmer) | 28 Aug 2018 |
DAY07 - post-M264 BS4 LuBot via Ekspla | |||
7.01 | laser data test | LuBot laser test data set @ 410 nm | 28 Aug 2018 |
DAY08 - post-M264 BS4 LuBot via Ekspla at 410 nm saturated and unsaturated | |||
8.01 | laser data test | LuBot laser test satu and unsat data set @ 410 nm | 28 Aug 2018 |
Compilation of day data | |||
20.01 | UV nets->BS link | RS pow lix data are negative- We found some nevative values in the low pixel on H19-01 and are checking we see them here | 31 Jan 2019 |
Collated data sets | |||
100.01 | Track Movement | Track movement of broad band sources | 28 Aug 2018 |
DAY 1 & 2 EMAIL from Mike |
Hi Steph! |
DAY 3 EMAIL from Mike |
Hi Steph, |
DAY 4 EMAIL from Mike |
Hi Steph, |
DAY 5 EMAIL from Mike |
Hi Steph, |
EMAIL from Mike about Es rsp |
I get the same same: <attached: ob64bs04Es01.png> Mikes and My math for Es system response matches up. So that is good. But does not explain why the postcal response is not MUCH higher.
|
DAY 6 EMAIL from Mike |
Hi Steph, |
DAY 7 EMAIL from Mike |
Hi Steph,
|
DAY 8 EMAIL from Mike |
Hi Steph, The BS4 was removed from MOBY last Thursday, 19-Jul-2018 - I don't know if you heard but there will be no BS nor RS on MOBY266. The camera in the BS4 has a flat window, so this will be going back to Andor for mods. On Monday = GMT 17-Jul I got a set of Ekspla laser saturation data via post-M264 BS04cfg005 LuBot = post-L271 day08. These are the last post-M264 BS data. Most curious I am what you can do with these! Only one laser line was scanned @ 410 nm = same as day07. The laser power was set at max and was not changed. BS scans were @ 2,4,6 sec = not saturated, and @ 8,10,15,20,40,80 sec = saturated. All scan sets were 2dk,4lt,2dk. It looks to I that the 40 & 80 sec scans bloomed out of the track, so I'm thinking these will not be of much use looking at the wings (?). I'm hoping we can figure out how far I can go with collecting saturated data such that it's not just junk! I spent too much time getting a TEC'ed Si Photodiode monitor on the sphere. Carol has been calibrating the "primary" SiPD so I set up our "secondary" kine. The PD is a Gigahertz TD1101-1. It's ampere output went to a brand new pre-amp = Gentec SDX-1226-2. The per-amp got ±15 Vdc power from a HP 6237B power supply. The Voltage output of the per-amp was measured via our SLM setup = 3497A DVM + Labview software w/ chan #1 cfg setup as for shunt monitor during Lu & Ed cals = high resolution 6.5 digit mode. I'm unsure if you've seen the output from this LabView Vdc6_5.vi before? I did forget to log a thermistor for ambient temperature as I usually do via Vdc6_5.vi + 34970A, so, unfortunately that is missing from the PD log - sorry aboot that. I think it will be instructive to get this SiPD logging figured out, since that will be part of our new SLC data routine... I think I'll write a better .vi for PD logging - one issue is that the pre-amp gain will likely need to be adjusted depending on the laser power output, and on this pre-amp the gain is set manually via a dial on the box, so the operator will need to update the gain setting in the logging software. There are 36x pairs of .fits files @ /ftp1/Mike/L271/bs4cal/day08/ , plus 2x logs = *_putty.log & *_pdmon.txt. There is 1x new log sheet for day08 @ /ftp1/Mike/L271/bs4cal/doc/ . There are 8x photos of the PD setup @ /ftp1/Mike/L271/bs4cal/photos/ . Many Mahalo, MF
|